OpenStudy (cherie_magee):

Solve the following system of equations. 2x + y = 3 x = 2y - 1

OpenStudy (anonymous):

2(2y-1)+y=3 4y-2+y=3 5y=5 y=1 x=2(1)-1 x=1

OpenStudy (anonymous):

First, change the second equation to match the first. You should end up with x-2y=-1 Now, multiply it by -2 to get the following: -2x+4y=2 This allows you to eliminate the x in both equations. Add the ys and the constants, then divide by the coefficient of y to get the y value. Then, take the value and plug it in to either equation to get x.